5. What is the probability that a market inspector will discover, greater than five violations of the public health code?





1
Expert's answer
2022-03-01T09:17:29-0500

Let the random variable "X" represent the number of violations of the public health code.

The probability that a market inspector will discover greater than 5 violations of health code is

"P(X\gt5)=1-P(X\le5)=1-P(X=0)"


"-P(X=1)-P(X=2)-P(X=3)"

"-P(X=4)-P(X=5)"

"=1-0.41-0.22-0.17-0.13-0.05-0.02"

"=0"